OpenStudy (vera_ewing):

If a football player passes a football from 4 feet off the ground with an initial velocity of 36 feet per second, how long will it take the football to hit the ground? Use the equation h = -16t^2 + 36t + 4.

OpenStudy (vera_ewing):

@dtan5457 This is the last one!

OpenStudy (vera_ewing):

@iambatman please please please help me? This is my last problem!

OpenStudy (vera_ewing):

I'm pretty sure it's A but idk....

OpenStudy (anonymous):

Can you show me your work?

OpenStudy (anonymous):

You just have to set it = 0 and solve

OpenStudy (vera_ewing):

Ugh it's on paper... will you wait for me to take a picture?

OpenStudy (dtan5457):

Quadratic formula again..lol.

OpenStudy (anonymous):

Yes, take your time, and yes quadratic formula is good!

OpenStudy (vera_ewing):

@dtan5457 it's A right?

OpenStudy (dtan5457):

-16t^2 + 36t + 4=0 i'd start by diving this by 4.

OpenStudy (vera_ewing):

-4t^2+9t+1=0/4

OpenStudy (dtan5457):

no need to divide the 0

OpenStudy (vera_ewing):

oh ok now what?

OpenStudy (dtan5457):

quadratic formula?

OpenStudy (anonymous):

Complete the square

OpenStudy (anonymous):

Yes, you can try using the quadratic formula as well

OpenStudy (vera_ewing):

t=[ − (−9)±√{(−9)²− 4×4×(− 1)}]/(2×4) =[ 9 ±√{81+ 16}]/(8) =[ 9±√{97}]/(8) =2. 356 Seconds

OpenStudy (vera_ewing):

but that's not one of my answer choices that's why i'm stuck :/

OpenStudy (anonymous):

You sure?

OpenStudy (dtan5457):

you don't have to simplify to actual seconds here

OpenStudy (anonymous):

Check your options again.

OpenStudy (dtan5457):

The answer is in there.

OpenStudy (vera_ewing):

Ohhhh wait hahaha

OpenStudy (dtan5457):

once again you did the b^2-4ac wrong

OpenStudy (dtan5457):

it's not a

OpenStudy (vera_ewing):

So is it C or D?

OpenStudy (dtan5457):

D is not even in the correct form

OpenStudy (vera_ewing):

Oh ok. Thanks both of you @iambatman @dtan5457 :)

OpenStudy (dtan5457):

No problem. any more questions?

OpenStudy (vera_ewing):

Nope that was the last one :) thank goodness lol

OpenStudy (dtan5457):

;p do things earlier next time :DD